Physics test review problem

You've got a flat tire. To lift your car, you make a homemade lever. A very light 1.6--long handle part is pushed down on the right side of the fulcrum and a 0.050-m

long part on the left side supports the back of the car.


How hard must you push down on the handle so that the lever exerts an 9700- force to lift the back of the car?
Answer #1

Considering equilibrium of moment

Fd = fr

F = 9700 * 0.05/1.6

F = 303.125 N


answered by: Allen
Answer #2

SOLUTION :


Let force to be applied on the handle be F (Newtons).


Moment about fulcrum :


9700 * 0.05 = F * 1.6 

=> F = 9700 * 0.05 / 1.6

=> F = 303.125 N (ANSWER).

answered by: Tulsiram Garg